OpenStudy (anonymous):

derivative of 5x sin(2x)

OpenStudy (anonymous):

@Mertsj

OpenStudy (kc_kennylau):

chain rule+product rule :D

OpenStudy (anonymous):

could you solve this one since it has trig functions plzz

OpenStudy (kc_kennylau):

for product rule, f=5x and g=sin(2x)

OpenStudy (kc_kennylau):

for chain rule, y=sin(2x) and u=2x

OpenStudy (anonymous):

hence what is the derivative?

OpenStudy (kc_kennylau):

\[\frac d{dx}\sin(2x)=\cos(2x)\cdot\frac d{dx}2x=2\cos(2x)\]

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

you forgot about 5x

OpenStudy (kc_kennylau):

i'm just doing the part with trigonometric function

OpenStudy (kc_kennylau):

you have to apply the product rule yourself xP

OpenStudy (anonymous):

k

OpenStudy (kc_kennylau):

b/c this website ain't a homework-doing website xP

OpenStudy (anonymous):

lol

OpenStudy (kc_kennylau):

think of anything?

OpenStudy (anonymous):

yah 5(2)cos(2x)

OpenStudy (kc_kennylau):

do you recall the product rule? what's the product rule? :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

hey can u help me with another question plz? just tell me which one lol, I dont know hwo to answer it

OpenStudy (anonymous):

The graph of y = x6 + 4.5x5 − 10x4 + 5x − 10 is concave upward for all values of x such that: (Points : 4) 0 < x < 4 −4 < x < 1 all real x x < −4 or x > 1 x < 0 or x > 4

OpenStudy (kc_kennylau):

\[\frac d{dx}f(x)g(x)=f'(x)g(x)+g'(x)f(x)\]

OpenStudy (kc_kennylau):

and you werent correct

OpenStudy (anonymous):

The graph of y = x6 + 4.5x5 − 10x4 + 5x − 10 is concave upward for all values of x such that: (Points : 4) 0 < x < 4 −4 < x < 1 all real x x < −4 or x > 1 x < 0 or x > 4

OpenStudy (anonymous):

could u help me plz

OpenStudy (kc_kennylau):

\[\frac d{dx} 5x\sin(2x)=5\sin(2x)+10x\cos(2x)\]

OpenStudy (anonymous):

thnxs

OpenStudy (kc_kennylau):

https://www.desmos.com/calculator/u1aavwzeqz

OpenStudy (anonymous):

...

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

all real x?

OpenStudy (kc_kennylau):

lemme fink

OpenStudy (kc_kennylau):

find the second derivative of that function, if the second derivative is positive within that interval, it's concave up.

OpenStudy (anonymous):

could you do it for me plzz

OpenStudy (kc_kennylau):

no http://openstudy.com/code-of-conduct

OpenStudy (kc_kennylau):

Find its second derivative first

OpenStudy (anonymous):

5x Sin (2x).... 10xCos(2x)+5Sin(2x)